Grade 12th passPhysical Chemistry

Calculate the mass percentage of benzene (C6H6) and carbon tetrachloride (CCl4) if 22 g of benzene is dissolved in 122 g of carbon tetrachloride.

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Answer ::

Mass of solution = Mass of C6H6 + Mass of CCl4
= 22 g+122 g= 144 g
Mass % of benzene = 22/144 x 100 =15.28 %
Mass % of CCl4 = 122/144 x 100 = 84.72 %

Thanks